### v1

# The Dual-Route Model of Induction

Abstract
Prior work on in-context copying has shown the existence of induction heads, which attend to and promote individual tokens during copying. In this work we introduce a new type of induction head: concept-level induction heads, which copy entire lexical units instead of individual tokens. Concept induction heads learn to attend to the ends of multi-token words throughout training, working in parallel with token-level induction heads to copy meaningful text. We show that these heads are responsible for semantic tasks like word-level translation, whereas token induction heads are vital for tasks that can only be done verbatim, like copying nonsense tokens. These two “routes” operate independently: in fact, we show that ablation of token induction heads causes models to paraphrase where they would otherwise copy verbatim. In light of these findings, we argue that although token induction heads are vital for specific tasks, concept induction heads may be more broadly relevant for in-context learning.
1 Introduction
How do language models repeat sequences of tokens in-context? Imagine that you are asked to copy a random string of characters from a leaflet of paper into a notebook: ‘oane dnn t ephzawfeew eausr lthii’. This would be an achievable but laborious task, requiring careful attention to each subsequent character. Now, imagine instead that these characters are rearranged into the phrase ‘the false azure in the windowpane’. There are now two ways of copying the sequence: character-by-character, or by leveraging your understanding of English to copy large swaths of characters at a time.
LLMs can copy text using induction heads, a type of circuit found in decoder-only transformer models that enables them to copy sequences in-context (Elhage et al., 2021; Olsson et al., 2022). Given the sequence w|ax|wing...w, an induction head at the second occurrence of w would attend to and promote the earlier token ax by scanning for previous token information. Prior work has argued that induction heads are responsible for broader in-context learning capabilities (Olsson et al., 2022). However, it is unclear how attention heads operating on a token-by-token basis might handle “fuzzy” copying tasks like translation, which often requires conversion between words of differing token lengths (e.g., p|ommes| des| terre $\rightarrow$ pot|atoes). In this paper, we show that LLMs use concept induction heads in parallel with token induction heads to copy meaningful text.
Figure 1: The dual-route model of induction. LLMs develop token induction heads, which are used for verbatim copying, and concept induction heads, important for translation and “fuzzy” copying tasks. These two routes work in parallel to copy meaningful text.
The Dual-Route Model of Induction.
Psychologists who study reading in the brain describe two parallel routes through which people read: a sublexical route that converts letter strings into speech sounds, and a lexical route through which word meanings can be directly accessed as entire units (Marshall & Newcombe, 1966; 1973; Dehaene, 2009). If the sublexical route is damaged, patients exhibit a condition known as deep dyslexia, where they can understand the meanings of words without being able to access their sound or spelling: Marshall & Newcombe (1966) describe a patient who, after a brain injury, would read the word CANARY as “parrot” and COLLEGE as “school,” indicating that he could still access word meanings, but was unable to read individual graphemes.
In this work, we consider the possibility of an analogous dual-route model of induction in LLMs, where tokens are equivalent to graphemes. Given a piece of meaningful text, models can either copy by shifting individual subword tokens, or by accessing detokenized lexical information over an entire span of tokens. We characterize two types of induction:
Token Induction. Figure 1 illustrates token-level induction circuits discovered in previous work. Elhage et al. (2021) show that in two-layer transformers, models load previous token information into each hidden state, which allows induction heads in future layers to attend to these states and copy the corresponding token. We design a causal intervention to identify token induction heads in Section 2. When these heads are ablated, models lose the ability to copy sequences verbatim (Section 4), exhibiting “symptoms” analogous to patients with deep dyslexia (Hinton & Shallice, 1991; Zorzi et al., 1998).
Concept Induction. Figure 1 also depicts our current understanding of concept induction heads, which are responsible for copying lexical information. In this work, we focus on concepts that take the form of lexical items (multi-token words and entities), as discussed by Feucht et al. (2024). To isolate these heads, we define a concept copying score based on causal mediation for a simple multi-token copying task (Section 2). Instead of attending to the next token, we find that these heads attend to the ends of multi-token words (Section 3), where concept information is more likely to be stored (Meng et al., 2022; Geva et al., 2023; Nanda et al., 2023; Kaplan et al., 2025). Concept induction heads are vital for semantic copying tasks (Section 4), and output word representations that are language-agnostic (Section 5).
2 Token & Concept Copying Scores
2.1 Approach
To identify concept induction heads, we search for attention heads responsible for copying multi-token words by measuring the effects of causal interventions (Vig et al., 2020; Geiger et al., 2023). We hypothesize that if a head increases the probability of future tokens for multi-token concepts, it is actually copying the entire concept.
Figure 2 illustrates our approach. We first sample random tokens [^1] To avoid confounds from undertrained tokens (Land & Bartolo, 2024), we sample tokens by randomly selecting a document from the Pile (Gao et al., 2020), tokenizing it, and shuffling the token order. We adopt this approach for all random token sampling henceforth. to create an induction prompt $x_{1}x_{2}...x_{n}|x^{\prime}_{1}x^{\prime}_{2}...x^{\prime}_{n}$, using the newline token as a separator between the first and second repeating occurrences of $x_{1}...x_{n}$. Then, we append a single concept made of $m$ tokens $c_{1}...c_{m}$ to each half of the repeated sequence $s$. These concepts are sampled from a set of multi-token concepts $\mathcal{C}$, with lengths uniformly distributed over $2\leq m\leq 5$. Prompts are truncated so that the final token is always $c^{\prime}_{1}$ regardless of $m$. We set $n=30-m$.
This yields a clean prompt $p_{clean}=x_{1}x_{2}...x_{n}c_{1}...c_{m}|x^{\prime}_{1}x^{\prime}_{2}...x^{\prime}_{n}c^{\prime}_{1}$, which we corrupt by re-sampling all of the tokens in the first half: $p_{corrupt}=y_{1}y_{2}...y_{n+\ell}|x^{\prime}_{1}x^{\prime}_{2}...x^{\prime}_{n}c^{\prime}_{1}$. We patch the activations of each attention head $a^{(l,h)}$ (layer $l$, head $h$) from the penultimate token position of $p_{clean}$ (i.e., from $x^{\prime}_{n}$) into the same position in $p_{corrupt}$. Then, the concept copying score for head $(l,h)$ over concept set $\mathcal{C}$ is defined as
$\textsc{ConceptCopying}{(l,h)}=\frac{1}{|\mathcal{C}|}\sum_{c\in\mathcal{C}}{\left(P(c_{2}|a_{p_{clean}}^{(l,h)}\underset{x^{\prime}_{n}}{\rightarrow}p_{corrupt})-P(c_{2}|p_{corrupt})\right)}$ (1)
where $\underset{x^{\prime}_{n}}{\rightarrow}$ indicates that activations $a^{(l,h)}$ are being patched into the $p_{corrupt}$ context at the position corresponding to $x^{\prime}_{n}$. Because $c_{2}$ is predicted at the token position after $x^{\prime}_{n}$, we are measuring increase in probability for the “next-next token” (Pal et al., 2023; Wu et al., 2024). Our hypothesis is that if an attention head increases probability for the future token $c_{2}$, this may be because it is carrying information about the entire concept $c_{1}...c_{m}$.
We also want to find attention heads that are responsible for copying one token at a time (i.e., token induction heads). We run the same procedure to find these heads, with two differences: (1) We use random tokens $r_{1}...r_{m}\in\mathcal{R}$ instead of concepts, and (2) we measure the impact of each head on $P(r_{1})$, instead of $P(r_{2})$. This gives us our token copying score:
$\textsc{TokenCopying}{(l,h)}=\frac{1}{|\mathcal{R}|}\sum_{r\in\mathcal{R}}{\left(P(r_{1}|a_{p_{clean}}^{(l,h)}\underset{x^{\prime}_{n}}{\rightarrow}p_{corrupt})-P(r_{1}|p_{corrupt})\right)}$ (2)
For this experiment and for Section 3, we use the CounterFact dataset (Meng et al., 2022), which consists of subject-object relations (e.g., “Paul Chambers plays bass”). We sample a subset of subjects from this dataset to use as our set of concepts $\mathcal{C}$, where $|\mathcal{C}|=1024$. We could also use generic multi-token words, but results from Feucht et al. (2024) suggest that models treat both types of sequences similarly. We also measure scores for $P(c_{1})$ and $P(r_{2})$, but do not focus on them in this work (see Appendix A).
Figure 2: We patch the outputs of each attention head from bar in $p_{clean}$ to bar in $p_{corrupt}$ to see whether that head has an impact on the “next-next” token $P(c_{2})$, which is $P(\texttt{ax})$ in this example. Our hypothesis is that the heads that increase $P(\texttt{ax})$ in this setting actually carry the entire concept of “waxwing.” See Section 2.1 for notation details.
Model Hugging Face ID $|\mathcal{V}|$ $t$ Citation
Llama-2-7b meta-llama/Llama-2-7b-hf 32k 2T Touvron et al. (2023)
Llama-3-8b meta-llama/Meta-Llama-3-8B 128k 15T+ Grattafiori et al. (2024)
OLMo-2-7b allenai/OLMo-2-1124-7B 100k 4T OLMo et al. (2025)
Pythia-6.9b EleutherAI/pythia-6.9b 50k 0.3T Biderman et al. (2023)
Table 1: Models used in this paper. $|\mathcal{V}|$ is the model’s token vocabulary size, and $t$ is the number of tokens the model was trained on (in trillions).
2.2 Results
We calculate causal scores for all heads in each model [^2] We use the Hugging Face (Wolf et al., 2020) implementation of each model. in Table 1. This gives us two ways of ranking heads in a model: by concept copying score and by token copying score. Figure 3a shows that concept copier heads seem to be concentrated in mid-early layers, whereas token copier heads are more sporadic, and are more likely to appear at late layers. We find that there is little overlap between the top-$k$ heads of each ranking (Figure 30) and that token & concept scores are not correlated (Appendix C.2), implying that these are two separate roles.
As a baseline, we also calculate increases in probability for future random tokens $P(r_{2})$ when patching individual heads (Appendix A). This gives a sense of whether heads are copying several arbitrary tokens at a time. While some heads do increase the probability of future random tokens, their maximum intervention effects are at least three times smaller than heads that promote future entity tokens (for all models). We take this as preliminary evidence that concept copying heads copy meaningful units, not just arbitrary lists of tokens.
We also calculate copying scores throughout training checkpoints for OLMo-2-7b and Pythia-6.9b. Some of the top concept induction heads in OLMo-2-7b have high token copying scores before they develop into concept induction heads, suggesting that concept induction heads might develop from token induction heads. However, this pattern is not as clear for Pythia-6.9b. We show examples of head trajectories throughout training in Appendix B.
Figure 3: We use causal mediation to identify token copier heads that copy the next random token, as well as concept copier heads that copy future concept tokens. (a) Distribution of the top-16 token and concept copier heads across model layers. (b) Value-weighted attention scores over a repeating phrase for the top causally-ranked heads $(l.h)$ in Llama-2-7b. At the final token position “the”, token copier heads attend to the next token (“window.p.ane”), whereas concept copier heads attend to the end of the next word (“window.p.ane”).
3 Next-Token & Last-Token Attention Scores
Where do concept copying heads attend? In previous work, Olsson et al. (2022) found that token induction heads always attend to the next token—the one that they are about to copy. If concept copier heads transfer entire concepts at once, we would expect them to attend to the ends of multi-token words, where concept information is usually stored (Meng et al., 2022; Geva et al., 2023; Nanda et al., 2023). Figure 3b shows an example of attention patterns for both types of heads in Llama-2-7b. When the model is about to copy a multi-token word, its token copier heads attend to the next token (“window.p.ane”), while its concept copier heads attend to the end of the next word (“window.p.ane”).
3.1 Approach
To capture this attention behavior, we design a last-token matching score which measures how much attention is paid to the last token of a multi-token concept. First, we select a concept $c=c_{1}...c_{m}$ from CounterFact subjects $\mathcal{C}$, evenly sampling across concept lengths $2\leq m\leq 5$. We then construct random repeated sequences of tokens as before (see §2) and insert the concept on the left to create the prompt $x_{1}x_{2}...x_{n}c_{1}...c_{m}|x^{\prime}_{1}x^{\prime}_{2}...x^{\prime}_{n}$. The last-token matching score is then calculated as an average over the concept set $\mathcal{C}$:
[…]
where $\smash{A^{(l,h)}}$ represents the value-weighted attention scores for head index $h$ at layer $l$, and the square brackets indicate that we collect attention paid from $x^{\prime}_{n}$ to $c_{m}$.
For comparison, we also calculate attention paid to the next token over random sequences, known in previous work as “prefix-matching” scores (Olsson et al., 2022; Bansal et al., 2023). In this work, we refer to this as a head’s next-token matching score. Just like in §2, we use the same procedure that we do for last-token matching scores, with two differences: (1) we replace concepts with random spans of tokens $r=r_{1}...r_{m}$, and (2) we calculate attention paid to the next random token $r_{1}$, instead of the last token.
[…]
Following Kobayashi et al. (2020), all attention scores in this work are calculated using value-weighting; i.e., we multiply attention weights for each token by the $\smash{L^{2}}$ norm of their value vectors, then renormalize so the scores sum to one. This tends to account for cases where attention “rests” on unimportant tokens like <s> at the beginning of a prompt.
3.2 Results
Figure 4 shows next-token and last-token matching scores for heads of each type in Llama-2-7b, averaged over 2048 CounterFact entities. Token copier heads tend to have high next-token matching scores, whereas concept copier heads tend to have high last-token matching scores. [^3] Last-token attention scores are lower overall because attention can spread over the length of the concept; for example, “window.p” could be treated as a concept if the model guesses that the word is “windowpane” (we can already see this for some heads in Figure 3b). We show only the top 16 heads (ranked using the scores defined in Equations 1 and 2) for readability, but include top-64 scores for all models in Appendix C.1.
We also calculate correlations between causal and attention-based scores. As expected, token copying scores positively correlate with next-token matching scores (r=0.63, p$<$0.001) whereas concept copying scores correlate with last-token matching scores (r=0.44, p$<$0.001) for Llama-2-7b. We find significant correlations for all models; see Appendix C.2.
Some heads seem to attend to both next-tokens and last-tokens. In Llama-2-7b, head 11.2 is the third-highest token copier head, but also has a relatively high last-token copying score (and attends to ane in Figure 3b). We also see concept copier heads with next-token matching scores up to 0.20, such as head 13.23. This suggests that some heads may just copy the next “thing," whether that be the next token or the next concept.
Figure 4: Attention-based matching scores for the highest-scoring causal heads in Llama-2-7b. Consistent with prior work, heads that are responsible for copying the next random token have high next-token matching scores. On the other hand, heads that are responsible for promoting future entity tokens have the highest last-token matching scores.3
4 Lesioning Concept and Token Copier Heads
We have found a set of attention heads that attend to the ends of multi-token entities (Section 3) and help to copy the second tokens of those entities (Section 2). We hypothesize that these are concept induction heads, which copy by transferring entire concept representations at once. In this section, we show through targeted ablations that concept heads are vital for “fuzzy copying” tasks that deal with lexical semantics, whereas token induction heads are important for verbatim copying tasks.
4.1 Approach
First, we define a new “vocabulary list” task that requires models to copy a list of words in-context. Using word-pair data from Conneau et al. (2017) for five languages, we enumerate $n=10$ words, followed by a parallel list of English translations for each word (see Appendix D.2 for full examples). Using data from Nguyen et al. (2017), we also build similar prompts for uppercased and title-cased English words, synonyms, and antonyms. We calculate first-token accuracy and exclude word pairs that have the same first token. Finally, we add two verbatim tasks, where the first and second list are identical: English copying, using words from Conneau et al. (2017), and “nonsense copying,” using randomly-sampled tokens. Models are evaluated on 1024 prompts per task.
For these tasks, we mean-ablate (Wang et al., 2023a) sets of concept and token induction heads to observe impact on model performance. We use causal scores from Section 2 to rank heads in each model twice: once according to concept copying score, and once according to token copying score. We ablate the top-$k$ heads in each of these rankings by replacing their activations across all token positions with their mean activations on random Pile documents ($n=1024$).
4.2 Results
Figure 5 shows ablation results for Llama-2-7b. We see two separate “routes” through which models can copy: a semantic route, via concept induction heads (heads with high concept copying scores), and a verbatim route, using token induction heads. When concept induction heads are ablated, we see a drop in translation, synonym, and antonym accuracy, with no effect on surface-level tasks. When token induction heads are ablated, nonsense copying fails, but the model can still use concept heads to complete the rest of the tasks. English copying remains unaffected for both types of ablation (e.g. pea|coat $\rightarrow$ pea|coat), as it can be solved using either type of induction. Like English copying, uppercasing tasks can also be done semantically or without regard to meaning. We report similar results for other models in Appendix D. For all models, ablation of concept induction heads damages semantic tasks much more than it does surface-level tasks. Llama-3-8b and OLMo-2-7b behave similarly to Llama-2-7b, but token ablation results for Pythia-6.9b are less clear-cut.
In Appendix D.3, we repeat this experiment while controlling for word length to test whether we are simply distinguishing between multi- and single-token tasks. For single-token words, we see the same pattern, with a weaker (but still distinct) separation between translation and verbatim copying when ablating concept heads. Thus, in cases where words are constrained to a single token, token heads may also be able to assist in semantic tasks.
Figure 5: Mean-ablating the top-$k$ copier heads for “vocabulary list” tasks in Llama-2-7b. Ablating token copier heads damages copying for nonsense tokens, without affecting tasks that can be performed semantically. Ablating concept copier heads damages performance for translation, synonyms, and antonyms, without affecting surface-level copying. English copying, which can be done either semantically or token-by-token, remains high for both types of ablation, as do uppercasing tasks (which can be done without regard to semantics). We plot results relative to models’ original task accuracies; see Appendix D for details.
Qualitative Examples.
To get a sense of how model outputs look when token induction heads are ablated, we prompt token-ablated models to copy entire sentences verbatim. Ablating a small number $k$ of token attention heads results in “paraphrasing” behavior instead of copying behavior; in other words, the model is able to copy the meaning of the sentence, but doesn’t get every token exactly right. Box 4.2 shows that Llama-2-7B starts to paraphrase when $k=32$ token induction heads are ablated. [^4] To encourage copying, we feed the sequence twice; we omit the first occurrence here for brevity. We find that $k=32$ is a sweet spot that allows us to observe this effect without causing the model to stop copying altogether. Box 4.2 shows ablated Llama-3-8b generations when prompted to copy a piece of Python code. When token induction heads are ablated, the model still copies the meaning of the original snippet, but writes it using list comprehension. We provide examples of this paraphrasing behavior for all models in Appendix D.4.
Box 4.1: (Llama-2-7b) Original Model vs. Top-32 Token Induction Heads Ablated I have reread, not without pleasure, my comments to his lines, and in many cases have caught myself borrowing a kind of opalescent light from my poet’s fiery orb. I have reread, not without pleasure, my comments to his lines, and in many cases have caught myself borrowing a kind of opalescent light from my poet’s fiery orb. I have reread my comments on his lines, and I have caught myself many times borrowing from his fiery orb a kind of opalescent light.
Box 4.2: (Llama-3-8b) Original Model vs. Top-32 Token Induction Heads Ablated foo = [] for i in range(len(bar)):     if i % 2 == 0:         foo.append(bar[i]) foo = [] for i in range(len(bar)):     if i % 2 == 0:         foo.append(bar[i]) foo = [] for i in range(len(bar)):     if i % 2 == 0:         foo.append(bar[i]) foo = [bar[i] for i in range( len(bar)) if i % 2 == 0]
5 Concept Induction is Language-Agnostic
If concept induction heads are important for translation tasks, what do they output? We hypothesize that the representations being copied by concept induction heads are abstract representations of word meanings. In other words, we posit that concept induction heads have the same activations when copying “waxwing” as they do when copying “свиристель,” as these two words refer to the same concept, and are only expressed differently on a surface level. This is in line with previous work suggesting that concept information in LLMs may be language-agnostic (Wendler et al., 2024; Dumas et al., 2024; Brinkmann et al., 2025), though Schut et al. (2025) argues concepts for some tasks may be biased towards English.
5.1 Approach
We adopt a similar experimental approach to Dumas et al. (2024). They provide a dataset of word-level translation prompts, where the model is shown five $\smash{(\ell^{(in)},\ell^{(out)})}$ pairs and prompted to translate a word $w$ from $\smash{\ell^{(in)}\rightarrow\ell^{(out)}}$. Following their approach, we define a source prompt $\smash{s:\ell^{(in)}_{s}\rightarrow\ell^{(out)}_{s}}$ and a base prompt $\smash{b:\ell_{b}^{(in)}\rightarrow\ell^{(out)}_{b}}$, where input and output languages differ between source and base prompts. For example, a source prompt might translate from Spanish to Italian, and a base prompt might translate from Japanese to Chinese. These prompts translate two different words $w_{s}$ and $w_{b}$. Their target outputs are those words expressed in their respective output languages: $\smash{w_{s}:\ell_{s}^{(out)}}$ and $\smash{w_{b}:\ell^{(out)}_{b}}$.
We patch the set of activations $a_{s}^{k}=\{a_{s}^{(l,h)}|(l,h)\in C^{k}\}$ of the top-$k$ concept copying heads $\smash{C^{k}}$ from the last token position of $s$ into the last token position of $b$. We generate $t=10$ tokens in this manner with NNsight multi-token generation (Fiotto-Kaufman et al., 2025), intervening for each newly-generated token. Then, we measure performance by evaluating accuracy of model generations. Specifically, given a generated string $w$, we consider $w$ equal to a ground truth label if it is contained within or contains the ground truth string. Dumas et al. (2024) provide multiple possible translations for output words, and $w$ is marked correct if it is equal to any of these synonyms.
5.2 Results
Figure 6 shows results for Llama-2-7b when patching outputs from a Spanish-Italian prompt into a Japanese-Chinese prompt. Patching concept induction heads causes the model to output the source word $w_{s}$ in the base language $\smash{\ell^{(out)}_{b}}$. For example, patching the outputs of concept heads from a Spanish-Italian prompt translating “child” into a Japanese-Chinese prompt causes the model to output the word for “child” in Chinese. This intervention is the most effective at about $k=80$, which also corresponds to the point where translation and nonsense copying are the most separate in Figure 5c. We show similar results for Llama-3-8b and more language pairs in Appendix E. These experiments suggest that concept induction heads transport semantic representations that are expressible in multiple languages.
Figure 6: (a) Patching the top-$k$ concept induction head outputs from a Spanish-Italian prompt into a Japanese-Chinese prompt. (b) Patching concept induction heads changes the concept output by the model, without affecting the language (i.e. the Spanish word is output in Chinese instead of Italian). This effect is strongest for $k=80$, which is also where the largest separation between semantic and literal copying performance is in Figure 5c.
6 Function Vector Heads Versus Concept Induction Heads
Figure 7: (a) Patching concept heads changes the output meaning without affecting language (Figure 6), while patching FV heads changes the output language without affecting meaning. (b) Patching FV heads from a French-English prompt into a German-Russian prompt ($n=128$) flips the output language to English (green) without changing the concept (red).
Function vector (FV) heads are attention heads whose outputs help to promote in-context tasks; e.g., the outputs of FV heads for a few-shot antonym task can also cause the model to output antonyms in new contexts (Todd et al., 2024). We notice that the highest-scoring concept copier head is also the highest-scoring FV head by a large margin for both Llama models. We calculate correlations between FV scores and concept copying scores and find significant, albeit weak, positive correlations for all models, suggesting there may be some relationship between these two types of heads (Figure 38).
However, while FV and concept induction heads seem to be related, they seem to play two distinct roles in semantic copying tasks. We focus on translation, and patch outputs of FV heads between translation prompts using the same approach that we do for concept induction heads in Section 5. Our experimental setup and data is identical, except we patch activations $\smash{a_{s}^{k}=\{a_{s}^{(l,h)}|(l,h)\in F^{k}\}}$, where $\smash{F^{k}}$ represents the top-$k$ FV heads for a given model. We then measure the output of the base word in the source output language, $\smash{w_{b}:\ell_{s}^{(out)}}$. Figure 7 shows these results for Llama-2-7b. For the exact same example, FV heads cause the model to output the same concept in a different language, whereas concept induction heads cause the model to output a different concept in the same language.
We also replicate ablations from Section 4 using FV rankings instead of concept and token copying rankings (Appendix F). Ablation of FV heads damages performance for non-verbatim tasks significantly, which makes sense: models cannot perform a task without knowing what the task is.
7 Related Work
Induction Heads and ICL. The in-context learning capabilities of LLMs as demonstrated by Brown et al. (2020) motivate a body of research on ICL (Dong et al., 2024; Lampinen et al., 2024). We build on the discovery of Elhage et al. (2021) and Olsson et al. (2022), which characterizes token induction heads. Concurrent with our work, Yin & Steinhardt (2025) argue that function vector (FV) heads (Todd et al., 2024), not token induction heads, are primarily responsible for few-shot ICL performance. Similarly, Yang et al. (2025) find that FV heads, which they call symbolic induction heads, perform induction over abstract variables. Akyürek et al. (2024) document n-gram heads, an n-gram generalization of induction head behavior, and Ren et al. (2024) study semantic induction heads that encode syntactic and semantic relations. Our work differs from these previous studies in that we identify concept induction heads via their ability to copy multi-token concepts.
Concept Representations in LLMs. We build upon previous work showing that LLMs contain internal representations of “concepts” beyond individual tokens (Kaplan et al., 2025; Hewitt et al., 2025). Some work has studied how individual tokens are converted into abstract concept representations via neurons (Elhage et al., 2022; Gurnee et al., 2023; Nanda et al., 2023) or attention heads (Correia et al., 2019; Ferrando & Voita, 2024). Other work has shown that this concept-level information is stored at the ends of multi-token entities/phrases across various settings for factual recall (Meng et al., 2022; Hernandez et al., 2024; Feucht et al., 2024; Ghandeharioun et al., 2024; Ferrando et al., 2025), and classification tasks (Wang et al., 2023b; Tigges et al., 2024; Marks & Tegmark, 2024). Our work identifies particular heads that transport this conceptual information between token positions.
8 Conclusion
In this work we investigate how LLMs copy lexical information. We find concept induction heads that are responsible for copying multi-token words, and whose representations are expressible across languages. These heads act as a second “route” through which models can copy meaningful text, alongside previously discovered token induction heads (Olsson et al., 2022). Concept induction heads are one example of how LLMs might use induction in a general way to transport abstract contextual information between hidden representations.
Acknowledgments
Thanks to Chris Wendler, Clément Dumas, and Kerem Sahin for helpful discussions on multilingual representations. We thank Hiba Ahsan and Can Rager for comments on our draft. SF also thanks Adrian Chang, Jerry H. Wang, and Emily Wang for feedback throughout the process of writing this paper. We thank the Center for AI Safety for providing computational resources used for certain experiments. Experiments are implemented in NNsight (Fiotto-Kaufman et al., 2025). SF, ET, and DB are funded by a grant from Open Philanthropy.
Ethics Statement
In this work, we study models trained primarily on English text. Although we claim that these models have representations that are agnostic to language, we acknowledge that these representations are likely biased towards English. In other words, they may disproportionately reflect the perspectives of people from English-speaking regions.
Appendix A Token & Concept Copying Scores
We provide copying scores for Llama-2-7b in Figure 9 and Figure 10, for Llama-3-8b in Figure 11 and Figure 12, for OLMo-2-7b in Figure 13 and Figure 14, and for Pythia-6.9b in Figure 15 and Figure 16.
Although we calculate intervention scores for the next token over entities $(c_{1})$ and the next-next token for random tokens $(r_{2})$, we do not focus on these scores in our work. In the former case, heads that are responsible for promoting the next concept token (i.e., $c_{1}$) could be either concept induction heads or token induction heads; we assume that this score would not help us to differentiate between the two. The latter score gives us information on heads that copy multiple arbitrary tokens (e.g., $r_{2}$) at a time (heads that copy “lists” of random tokens). Careful comparison of the right-hand plots for entity tokens versus random tokens (e.g., Figure 9 vs. Figure 10) shows that these heads have weaker effects, which helps to support our hypothesis that concept heads are not just “copying lists of tokens.” However, we do not utilize these scores further in this work. In Figure 8, we plot the concept copying score against token copying score for each model. We find that concept and token copying scores are either negatively correlated (Llama-2-7b, OLMo-2-7b, and Pythia-6.9b) or uncorrelated (Llama-3-8b). We take this to mean that concept copying and token copying are two distinct roles.
Figure 8: Concept copying scores plotted against token copying scores for every model. These scores are either not correlated or negatively correlated. In other words, There are few heads, if any, that are causally important for both random next-token copying and copying of future concept tokens.
Figure 9: Probability differences when patching head activations for Llama-2-7b over entity tokens.* The right-hand side shows concept copying scores; we do not utilize the left-hand scores in this work. We patch at token position $x^{\prime}_{n}$ and measure increase in probability for $c_{1}$ at that token position (left), and increase in probability for $c_{2}$ at the following token position. See Figure 2 and Section 2 for details.
*Head 14.1 scores an order of magnitude higher than the second-highest scoring head on the right plot. Its concept copying score is 0.0011. For visibility, we scale instead by the second-highest concept copying score.
Figure 10: Probability differences when patching head activations for Llama-2-7b over random tokens. The left-hand side shows token copying scores; we do not utilize the right-hand scores in this work. We patch at token position $x^{\prime}_{n}$ and measure increase in probability for $r_{1}$ at that token position (left), and increase in probability for $r_{2}$ at the following token position. See Figure 2 and Section 2 for details.
Figure 11: Probability differences when patching head activations for Llama-3-8b over entity tokens. The right-hand side shows concept copying scores; we do not utilize the left-hand scores in this work. We patch at token position $x^{\prime}_{n}$ and measure increase in probability for $c_{1}$ at that token position (left), and increase in probability for $c_{2}$ at the following token position. See Figure 2 and Section 2 for details.
Figure 12: Probability differences when patching head activations for Llama-3-8b over random tokens. The left-hand side shows token copying scores; we do not utilize the right-hand scores in this work. We patch at token position $x^{\prime}_{n}$ and measure increase in probability for $r_{1}$ at that token position (left), and increase in probability for $r_{2}$ at the following token position. See Figure 2 and Section 2 for details.
Figure 13: Probability differences when patching head activations for OLMo-2-7b over entity tokens. The right-hand side shows concept copying scores; we do not utilize the left-hand scores in this work. We patch at token position $x^{\prime}_{n}$ and measure increase in probability for $c_{1}$ at that token position (left), and increase in probability for $c_{2}$ at the following token position. See Figure 2 and Section 2 for details.
Figure 14: Probability differences when patching head activations for OLMo-2-7b over random tokens. The left-hand side shows token copying scores; we do not utilize the right-hand scores in this work. We patch at token position $x^{\prime}_{n}$ and measure increase in probability for $r_{1}$ at that token position (left), and increase in probability for $r_{2}$ at the following token position. See Figure 2 and Section 2 for details.
Figure 15: Probability differences when patching head activations for Pythia-6.9b over entity tokens. The right-hand side shows concept copying scores; we do not utilize the left-hand scores in this work. We patch at token position $x^{\prime}_{n}$ and measure increase in probability for $c_{1}$ at that token position (left), and increase in probability for $c_{2}$ at the following token position. See Figure 2 and Section 2 for details.
Figure 16: Probability differences when patching head activations for Pythia-6.9b over random tokens. The left-hand side shows token copying scores; we do not utilize the right-hand scores in this work. We patch at token position $x^{\prime}_{n}$ and measure increase in probability for $r_{1}$ at that token position (left), and increase in probability for $r_{2}$ at the following token position. See Figure 2 and Section 2 for details.
Appendix B Token and Concept Induction Heads Throughout Training
We include analysis of causal and attention-based induction scores over time for OLMo-2-7b and Pythia-6.9b (models that provide access to training checkpoints). In Figures 17 and 19, we show token and concept copying scores from Section 2 over checkpoints. Figures 18 and 20 show next-token and last-token attention scores from Section 3 over checkpoints.
Figure 17: Causal copying scores from Section 2 for OLMo-2-7b throughout training checkpoints. Top row: top-4 token induction heads. Bottom row: top-4 concept induction heads.
Figure 18: Attention-based matching scores from Section 3 for OLMo-2-7b throughout training checkpoints. Top row: top-4 token induction heads. Bottom row: top-4 concept induction heads.
Figure 19: Causal copying scores from Section 2 for Pythia-6.9b throughout training checkpoints. Top row: top-4 token induction heads. Bottom row: top-4 concept induction heads.
Figure 20: Attention-based matching scores from Section 3 for Pythia-6.9b throughout training checkpoints. Top row: top-4 token induction heads. Bottom row: top-4 concept induction heads.
Appendix C Next-Token and Last-Token Matching Scores
C.1 Attention Scores for Top 64 Heads
Figure 21: Next-token and last-token matching scores from Section 3 for the top-64 Llama-2-7b token copier heads. Heads that are also in the top-64 concept heads are bolded.
Figure 22: Next-token and last-token matching scores from Section 3 for the top-64 Llama-2-7b concept copier heads. Heads that are also in the top-64 token heads are bolded.
Figure 23: Next-token and last-token matching scores from Section 3 for the top-64 Llama-3-8b token copier heads. Heads that are also in the top-64 concept heads are bolded.
Figure 24: Next-token and last-token matching scores from Section 3 for the top-64 Llama-3-8b concept copier heads. Heads that are also in the top-64 token heads are bolded.
Figure 25: Next-token and last-token matching scores from Section 3 for the top-64 OLMo-2-7b token copier heads. Heads that are also in the top-64 concept heads are bolded.
Figure 26: Next-token and last-token matching scores from Section 3 for the top-64 OLMo-2-7b concept copier heads. Heads that are also in the top-64 token heads are bolded.
Figure 27: Next-token and last-token matching scores from Section 3 for the top-64 Pythia-6.9b token copier heads. Heads that are also in the top-64 concept heads are bolded.
Figure 28: Next-token and last-token matching scores from Section 3 for the top-64 Pythia-6.9b concept copier heads. Heads that are also in the top-64 token heads are bolded.
We include expanded versions of Figure 4 for the top-64 token and concept copier heads for each model. See Figures 21 through 28. We see that concept copier heads tend to have high last-token matching scores (§3), but also that each model has a few heads that seem to do both token and concept induction. Heads that appear in the top-64 of both token-copying scores and concept-copying scores are bolded.
C.2 Correlations With Causal Scores
Figure 29 shows correlations between causal scores (Section 2) and attention-based scores (Section 3). For token induction, we compare token copying scores with next-token matching scores. For concept induction, we compare concept copying scores with last-token matching scores. Although these scatterplots appear quite noisy, there does seem to be a significant correlation between attention paid to the next/last token and propensity to copy tokens/entities respectively. Correlations for concept induction are comparable to correlations for token induction heads (which have been established in prior work).
Figure 29: Relationship between attention-based matching scores (Section 3) and causal copying scores (Section 2). Top: correlations between next-token matching scores and token copying scores for each head, indicators of token-level induction. Bottom: correlations between last-token matching scores and concept copying scores for each head, indicating concept induction.
Figure 30: Number of overlapping token and concept copying heads for increasing values of $k$. For smaller values of $k$, OLMo-2-7b and Pythia-6.9b have more overlapping heads than Llama models.
Appendix D Lesioning Concept and Token Copier Heads
D.1 Full Results
We show the results of ablating token copier heads and concept copier heads on our “vocabulary list" tasks for all four models in Figures 31-34. We also show the number of heads that overlap for increasing choices of $k$ in Figure 30. Mean-ablation is done using average activations of all heads over $n=1024$ Pile (Gao et al., 2020) documents.
Figure 31: Mean-ablating the top-$k$ copier heads for “vocabulary list” tasks in Llama-2-7b. As results are shown as a percentage of original task accuracy, we display these full model accuracies in the legend. Even though synonym accuracy is initially low, it is still unaffected by ablation of token induction heads.
Figure 32: Mean-ablating the top-$k$ copier heads for “vocabulary list” tasks in Llama-3-8b. As results are shown as a percentage of original task accuracy, we display these full model accuracies in the legend. In contrast to Llama-2-7b, capitalization tasks seem to use a blend of concept induction and token induction heads.
Figure 33: Mean-ablating the top-$k$ copier heads for “vocabulary list” tasks in OLMo-2-7b. As results are shown as a percentage of original task accuracy, we display these full model accuracies in the legend. Although English copying accuracy remains high in the right figure, ablation of concept heads in OLMo-2-7b does damage nonsense token accuracy; this indicates that there may be more overlap between these heads for OLMo-2-7b than for Llama models.
Figure 34: Mean-ablating the top-$k$ copier heads for “vocabulary list” tasks in Pythia-6.9b. As results are shown as a percentage of original task accuracy, we display these full model accuracies in the legend. Unlike previous models, semantic tasks (translation, synonyms, and antonyms) are also damaged when we ablate token induction heads (left). However, capitalization and English copying accuracy remains relatively high. We do not have a good explanation for this result. Because Pythia is trained on an order of magnitude fewer tokens than other models in this work (0.3 trillion, versus $\geq$ 2 trillion), it may be that concept induction and token induction are somewhat fused.
D.2 Examples of Prompts
We show examples of two semantic task prompts from Section 4 in Box D.2. We also show examples of two verbatim tasks in Box D.2.
Box D.1: Examples of Vocabulary List Prompts (French-English & Antonym) <s> French vocab:  1. completement  2. inconstitutionnalité  3. flocon  4. chiens  5. vaccinés  6. racontée  7. spécialités  8. parfumée  9. condensateur  10. saumons  English translation:  1. completely  2. unconstitutionality  3. flake  4. dogs  5. inoculated  6. told  7. specialties  8. fragrant  9. capacitor  10. <s> Words:  1. hydrous  2. enlarge  3. alternative  4. messy  5. traditionalism  6. simplicity  7. underwater  8. transitive  9. omit  10. contract  Antonyms:  1. anhydrous  2. cut  3. primary  4. neat  5. rationalism  6. difficulty  7. amphibious  8. intransitive  9. remember  10.
Box D.2: Examples of Vocabulary List Prompts (English & Nonsense Copying) <s> English vocab:  1. live  2. begin  3. stumble  4. good  5. ostracize  6. important  7. coin  8. colored  9. manner  10. recover  English vocab:  1. live  2. begin  3. stumble  4. good  5. ostracize  6. important  7. coin  8. colored  9. manner  10. <s> Vocab:  1. any insp  2. comes look  3. the like  4.points_  5.    fix  6. \$\$_  7.ence Camer  8. there  9. object currently  10. ercase  Vocab:  1. any insp  2. comes look  3. the like  4.points_  5.    fix  6. \$\$_  7.ence Camer  8. there  9. object currently  10.
D.3 Word Length Breakdown
We run the same ablation experiment as shown in Figure 5 for Llama-2-7b, except we control the number of tokens in each word. We run only on a subset of representative tasks and for $n=256$ prompts per task. Figure 35 shows these results. While single-token words still have a similar pattern to multi-token words, the effect is less drastic, likely because the difference between concepts and tokens is less clear for single-token words.
Figure 35: Ablations for Llama-2-7b, controlling the number of tokens in each word. When translating single-token French words to English, ablation of concept heads is less effective, suggesting that token heads may also do semantic copying when words are restricted to single tokens. However, the separation is still apparent, and is even clearer for two- and three-token words.
D.4 Qualitative Examples
We show examples of model generations with token induction heads ablated, expanding on examples shown in Section 4.2. We generate text with token induction heads mean-ablated at all token positions using the nnsight.LanguageModel.generate function (Fiotto-Kaufman et al., 2025), with default sampling. For the first three models, ablating $k=32$ token induction heads causes the model to start paraphrasing instead of directly copying, for natural language (Boxes D.4, D.4, D.4, D.4) and code snippets (Boxes D.4,  D.4,  D.4).
We notice that for $k=32$, Pythia-6.9b’s “paraphrases” are much poorer than other models (Box D.4, D.4, D.4). We scale back to $k=16$ and provide more than the first token of the copied sequence in Box D.4, finding that this causes Pythia-6.9b to copy more faithfully.
Interestingly, when paraphrasing a Python snippet, Llama-3-8b (Box 4.2) is the only model to write a snippet that is “correct,” i.e. semantically equivalent to the original input.
We also show what model outputs look like for vocabulary list prompts: Llama-2-7b generates sentences that use previously-seen words (Box D.4), Llama-3-8b copies the list but without numbering (Box D.4), and OLMo-2-7b (Box D.4) and Pythia-6.9b (Box D.4) fail to copy any of the previously-seen words.
Box D.3: (Llama-3-8b) Original Model vs. Top-32 Token Induction Heads Ablated I have reread, not without pleasure, my comments to his lines, and in many cases have caught myself borrowing a kind of opalescent light from my poet’s fiery orb I have reread, not without pleasure, my comments to his lines, and in many cases have caught myself borrowing a kind of opalescent light from my poet’s fiery orb I’ve reread my comments on his lines, and found that I’d borrowed from his fiery orb a kind of opalescent light
Box D.4: (OLMo-2-7b) Original Model vs. Top-32 Token Induction Heads Ablated I have reread, not without pleasure, my comments to his lines, and in many cases have caught myself borrowing a kind of opalescent light from my poet’s fiery orb I have reread, not without pleasure, my comments to his lines, and in many cases have caught myself borrowing a kind of opalescent light from my poet’s fiery orb I have read his lines again, and I have not found them without pleasure. In many cases, I have caught myself borrowing a kind of opalescent light from his fiery orb.
Box D.5: (Pythia-6.9b) Original Model vs. Top-32 Token Induction Heads Ablated I have reread, not without pleasure, my comments to his lines, and in many cases have caught myself borrowing a kind of opalescent light from my poet’s fiery orb I have reread, not without pleasure, my comments to his lines, and in many cases have caught myself borrowing a kind of opalescent light from my poet’s fiery orb I have reread my comments on his lines, and I have found them to be, in general, true, and I have found myself to be in agreement with them.
Box D.6: (Pythia-6.9b*) Original vs. Top-16* Token Induction Heads Ablated I have reread, not without pleasure, my comments to his lines, and in many cases have caught myself borrowing a kind of opalescent light from my poet’s fiery orb I have reread, not without pleasure, my comments to his lines, and in many cases have caught myself borrowing a kind of opalescent light from my poet’s fiery orb I have reread, not without pleasure, my comments to his lines, and have caught myself in many cases, borrowing a kind of opalescent light from my poet’s fiery orb
Box D.7: (Llama-2-7b) Original Model vs. Top-32 Token Induction Heads Ablated foo = [] for i in range(len(bar)):     if i % 2 == 0:         foo.append(bar[i]) foo = [] for i in range(len(bar)):     if i % 2 == 0:         foo.append(bar[i]) foo = [] for i in range(len(bar)):     if i % 2 == 0:         foo.append(bar[i]) foo = [item for sublist in     bar for item in sublist     if i % 2 == 0]
Box D.8: (OLMo-2-7b) Original Model vs. Top-32 Token Induction Heads Ablated foo = [] for i in range(len(bar)):     if i % 2 == 0:         foo.append(bar[i]) foo = [] for i in range(len(bar)):     if i % 2 == 0:         foo.append(bar[i]) foo = [] for i in range(len(bar)):     if i % 2 == 0:         foo.append(bar[i]) foo = [x for x in bar if     x % 2 == 0]
Box D.9: (Pythia-6.9b) Original Model vs. Top-32 Token Induction Heads Ablated foo = [] for i in range(len(bar)):     if i % 2 == 0:         foo.append(bar[i]) foo = [] for i in range(len(bar)):     if i % 2 == 0:         foo.append(bar[i]) foo = [] for i in range(len(bar)):     if i % 2 == 0:         foo.append(bar[i]) foo = foo.append(foo) print(foo)
Box D.10: (Llama-2-7b) Original Model vs. Top-32 Token Induction Heads Ablated English vocab: 1. hospital 2. visage 3. unlucky 4. lesbian 5. know 6. wizard 7. unfair 8. sound 9. cortical 10. obsolete English vocab:
1. hospital
2. visage
3. unlucky
4. lesbian
5. know
6. wizard
7. unfair
8. sound
9. cortical
10. obsolete
4. lesbian 5. know 6. wizard 7. unfair 8. sound 9. cortical 10. obsolete English vocabulary:
1. She was in hospital for a long time after the accident.
2. Her face was covered in bandages.
3. The accident was very unlucky.
4. She was a lesbian.
5. He knows everything about the wizard.
6. It is unfair to treat people like that.
7. The sound of the engine was very loud.
8. I don’t know why the doctor told me that I should eat more fish.
9. The new cortical implants can help the blind to see.
10. The
Box D.11: (Llama-3-8b) Original Model vs. Top-32 Token Induction Heads Ablated English vocab: 1. hospital 2. visage 3. unlucky 4. lesbian 5. know 6. wizard 7. unfair 8. sound 9. cortical 10. obsolete English vocab:
1. hospital
2. visage
3. unlucky
4. lesbian
5. know
6. wizard
7. unfair
8. sound
9. cortical
10. obsolete
English vocab: 1. hospital 2. visage 3. unlucky 4. lesbian 5. know 6. wizard 7. unfair 8. sound 9. cortical 10. obsolete English vocab:
hospital
visage
unlucky
lesbian
know
wizard
unfair
sound
cortical
obsolete
Box D.12: (OLMo-2-7b) Original Model vs. Top-32 Token Induction Heads Ablated English vocab: 1. hospital 2. visage 3. unlucky 4. lesbian 5. know 6. wizard 7. unfair 8. sound 9. cortical 10. obsolete English vocab:
1. hospital
2. visage
3. unlucky
4. lesbian
5. know
6. wizard
7. unfair
8. sound
9. cortical
10. obsolete
English vocab: 1. hospital 2. visage 3. unlucky 4. lesbian 5. know 6. wizard 7. unfair 8. sound 9. cortical 10. obsolete English to French
French to English
English to French
French to English
German to English
Russian to English
German to Russian
English to German<|endoftext|>
Box D.13: (Pythia-6.9b) Original Model vs. Top-32 Token Induction Heads Ablated English vocab: 1. hospital 2. visage 3. unlucky 4. lesbian 5. know 6. wizard 7. unfair 8. sound 9. cortical 10. obsolete English vocab:
1. hospital
2. visage
3. unlucky
4. lesbian
5. know
6. wizard
7. unfair
8. sound
9. cortical
10. obsolete
English vocab: 1. hospital 2. visage 3. unlucky 4. lesbian 5. know 6. wizard 7. unfair 8. sound 9. cortical 10. obsolete English vocabularies
[…]
Appendix E Concept Induction is Language-Agnostic
We show results from Section 5 for more languages, with both Llama models. Figure 36 shows results for the same experiment as in Figure 6, expanding evaluation to Llama-3-8b and two more language sets: patching from French-English into German-Russian, and patching from Russian-Spanish into English-Japanese.
Figure 36: Results for patching the top-$k$ concept induction heads from one translation prompt to another. Top-left is the same plot as Figure 6. We evaluate Llama-2-7b and Llama-3-8b on Spanish-Italian $\rightarrow$ Japanese-Chinese, French-English $\rightarrow$ German-Russian, and Russian-Spanish $\rightarrow$ English-Japanese. Interestingly, the language set for which concept patching is the least effective is the same language set for which FV head patching is most effective.
Appendix F Function Vector Versus Concept Induction
As mentioned in Section 6, we find weak correlations between FV and concept copying scores (Figure 38). We also find that ablation of FV heads also causes a large drop in performance for vocabulary list tasks that cannot otherwise rely on token induction heads (Figure 39). However, this does not necessarily mean that FV heads play the same role as concept induction heads. As Section 6 explains, patching FV heads and patching concept induction heads for the same prompt yields very different results. While patching FV heads changes the language that the model outputs (depending on the language pair), patching concept induction heads changes the meaning of the output word. In order for the model to do semantic copying tasks, it needs both FV heads and concept heads—therefore, if either are ablated, we see the same drop in translation accuracy. We show results for patching FV heads for Llama models for two language sets in Figure 37.
Figure 37: Results for patching the top-$k$ FV heads from one translation prompt to another. Top-left is the same plot as Figure 7. Patching FV heads from Spanish-Italian to Japanese-Chinese does not flip the output language to Italian, but patchiing from French-English to German-Russian flips the output language to English.
Figure 38: Correlations between FV scores and entity copying scores for all models. While we do find significant correlations, they are strongest for Llama models with strong outliers, and do not seem particularly strong for other heads and models.
Figure 39: Ablation experiment described in Section 4 where we ablate the top-$k$ function vector (FV) heads. We see that FV heads are also vital for semantic copying tasks. As Section 6 describes, this result is likely not due to an overlap with concept induction heads, but rather because FV heads help the model output the correct language (whereas concept heads copy the word meaning).